40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Sep 11 17:17:51 2012, Eric Q., Mike J., Configuration, Electronics, AS beam scan 
    Reply  Wed Sep 12 08:16:49 2012, Steve, Configuration, PEM, chamber must be sealed overnight! 
    Reply  Wed Sep 12 20:08:51 2012, ericq, Update, Electronics, AS beam scan 
       Reply  Thu Sep 13 19:59:43 2012, Mike J., Update, Electronics, AS beam scan 
          Reply  Fri Sep 14 01:35:55 2012, Mike J., Update, Electronics, AS beam scan PLOTS H_razor.jpegV_razor.jpeg
             Reply  Tue Sep 18 22:06:21 2012, Mike J., Update, Electronics, AS beam scan plots and chi-squared beam_width.pyH_raz.pngV_raz.png
Message ID: 7377     Entry time: Wed Sep 12 20:08:51 2012     In reply to: 7372     Reply to this: 7380
Author: ericq 
Type: Update 
Category: Electronics 
Subject: AS beam scan 

Quote:

We conducted a beam scan on the AP table of the AS beam. We used a lens to focus the beam onto a power meter, and slowly moved a razor blade across the beam using a micrometer, vertically and horizontally both in front of and behind the beam. We also had to block the beam next to the AS beam in order to do this, but is unblocked now. Mike will begin curve fitting the data to try and see if there is a different spot size given by the x-axis vs. the y-axis, and if the lens has any effect.

 [ericq, mikej, some input from zach]

After realigning the MC, the measurement was repeated this afternoon. This time, however, we isolated the beam from ITMY by misaligning ITMX. The beam looked somewhat elliptical to me, and Mike should have fits up tonight. Afterwards, ITMX was returned to the position I found it in, and the PMC shutter and access connector were closed. (Sorry about last night!)

ELOG V3.1.3-